# Forecasting Self-Similar User Traffic Demand Using Transformers in LEO Satellite Networks

### v1

@@ Abstract
Index Terms:
Low Earth Orbit (LEO), satellite, beam-hopping, forecasting, traffic, transformers, FARIMA
@@ I Introduction
In this work, we consider a Low Earth Orbit (LEO) satellite constellation where each satellite with a regenerative payload provides broadband services to multiple earth-fixed cells. In each cell, there can be a different number of users with varying traffic demands on the forward link (downlink). Given this, the aggregated demand at each cell would be different. A snapshot of this system, where a single LEO satellite serving six fixed-earth cells can be seen on the left side of Fig. 1. Considering this setting, different cells would experience varying, irregular traffic demands. Therefore, some cells may experience higher demand and become “hot-spots” whereas some other cells may remain “colder”. In order to compromise this need, the Beam-Hopping (BH) technology emerges as a solution to utilize limited radio resources more flexibly compared to fixed beams. With the BH technology, hot-spot or higher priority cells can be illuminated for a longer duration of time with higher hopping rates.
@@ II-B ARIMA
$Y_{t}$ can be modeled with Autoregressive Moving-Average (ARMA). An ARMA(p,q)(p,q) process is given by (12) where $\varepsilon_{t}$ is assumed to be i.i.d. with zero mean and finite variance $\sigma_{\varepsilon}^{2}$. $\phi(k)=1-\sum_{j=1}^{p}\phi_{j}k^{j}$ and $\psi(k)=\sum_{j=0}^{q}\psi_{j}z^{j}$ are polynomials with no common roots and all roots being outside the unit circle [13]:
@@ II-C FARIMA
For $|k|\leq 1$, $z\neq 1$ and $\Gamma(\cdot)$ being Gamma function:
@@ III Transformer based forecasting solution
where $d_{k}$ constant represents the dimension of $Q$ and $K$ [14]. Considering this, the $QK^{\top}$ operation has a quadratic run-time and memory complexity with respect to $d_{k}$ in vanilla transformers.
Zhou et al. [16] addressed this shortcoming by proposing a Prob-Sparse attention mechanism achieving $\mathcal{O}(d_{k}\log d_{k})$ for long sequence time-series forecasting. During the attention mechanism, instead of using all $d_{k}$, the proposed architecture uses only the top-$u$ queries where $u=c\ln{L_{Q}}$ and $c$ is a constant. They employ a max-mean criterion to select the top-$u$ queries, ultimately achieving $\mathcal{O}(d_{k}\log d_{k})$ run-time and space complexity. We consider this architecture, Informer, as a transformer-based solution.
@@ III-A2 Informer
Informer predicts a sequence at one forward operation rather than step-by-step. For a given Query and Key pair with each having a dimension of $L$, the run-time complexity of forecasting is $\mathcal{O}(L\log L)$ [16]. It is worth noting that the dimensions of Query and Key vectors depend on the length of the model dimension rather than the length of the input sequence.
@@ IV Simulation Results
TABLE I: Table for high demand traffic; $M=750$, $a=1.04$, $H=0.98$
@@
TABLE II: Table for high demand traffic; $M=500$, $a=1.6$, $H=0.7$
@@ V Conclusion
TABLE III: Table for high demand traffic; $M=250$, $a=1.9$, $H=0.55$

### v2

@@ I Introduction
In this work, we consider a LEO satellite constellation where each satellite with a regenerative payload provides broadband services to multiple earth-fixed cells. In each cell, there can be a different number of users with varying traffic demands on the forward link (downlink). Given this, the aggregated demand at each cell would be different. A snapshot of this system, where a single LEO satellite serving six fixed-earth cells can be seen on the left side of Fig. 1. Considering this setting, different cells would experience varying, irregular traffic demands. Therefore, some cells may experience higher demand and become “hot-spots” whereas some other cells may remain “colder”. In order to compromise this need, the Beam-Hopping (BH) technology emerges as a solution to utilize limited radio resources more flexibly compared to fixed beams. With the BH technology, hot-spot or higher priority cells can be illuminated for a longer duration of time with higher hopping rates.
@@ II-B ARIMA
$Y_{t}$ can be modeled with Autoregressive Moving-Average (ARMA). An ARMA(p,q)(p,q) process is given by (12) where $\varepsilon_{t}$ is assumed to be i.i.d. with zero mean and finite variance $\sigma_{\varepsilon}^{2}$. $\phi(k)=1-\sum_{j=1}^{p}\phi_{j}k^{j}$ and $\psi(k)=\sum_{j=0}^{q}\psi_{j}k^{j}$ are polynomials with no common roots and all roots being outside the unit circle [13]:
@@ II-C FARIMA
For $|k|\leq 1$, $k\neq 1$ and $\Gamma(\cdot)$ being Gamma function:
@@ III Transformer based forecasting solution
where $d_{k}$ constant represents the dimension of $Q$ and $K$ [14]. Considering this, the $QK^{\top}$ operation has a quadratic run-time and memory complexity in vanilla transformers.
Zhou et al. [16] addressed this shortcoming by proposing a Prob-Sparse attention mechanism achieving $\mathcal{O}(L\log L)$ for long sequence time-series forecasting. During the attention mechanism, instead of using all $L$, the proposed architecture uses only the top-$u$ queries where $u=c\ln{L_{Q}}$ and $c$ is a constant. They employ a max-mean criterion to select the top-$u$ queries, ultimately achieving $\mathcal{O}(L\log L)$ run-time and space complexity. We consider this architecture, Informer, as a transformer-based solution.
@@ III-A2 Informer
Informer predicts a sequence at one forward operation rather than step-by-step. For a given Query and Key pair with each having a dimension of $L$, the run-time complexity of forecasting is $\mathcal{O}(L\log L)$ [16]. It is worth noting that the dimensions of the Query and Key matrices depend not only on the input sequence length but also on the model dimension.
@@ IV Simulation Results
TABLE I: Table for the high demand traffic; $M=750$, $a=1.04$, $H=0.98$
@@
TABLE II: Table for the medium demand traffic; $M=500$, $a=1.6$, $H=0.7$
@@ V Conclusion
TABLE III: Table for the low demand traffic; $M=250$, $a=1.9$, $H=0.55$